### v1

# BonnBeetClouds3D: A Dataset Towards Point Cloud-based Organ-level Phenotyping of Sugar Beet Plants under Field Conditions

Abstract
Agricultural production is facing severe challenges in the next decades induced by climate change and the need for sustainability, reducing its impact on the environment. Advancement in field management through non-chemical weeding by robots in combination with monitoring of crops by autonomous unmanned aerial vehicles (UAVs) and breeding of novel and more resilient crop varieties are helpful to address these challenges. The analysis of plant traits is called phenotyping, and is an essential activity in plant breeding, it however involves a great amount of manual labor. With this paper, we address the problem of automatic fine-grained organ-level geometric analysis needed for precision phenotyping. However, the availability of real-world data for such fine-grained perception tasks in this domain is relatively scarce compared to other domains such as autonomous driving. To work towards closing this gap, we propose a novel dataset that was acquired using UAVs capturing high-resolution images of a real breeding trial. This has the big advantage of containing a multitude of plant varieties, leading to a great morphological and appearance diversity covered by our dataset. This enables the development of approaches for autonomous phenotyping that generalize well to different varieties. Based on overlapping high-resolution images from multiple viewing angles, we compute photogrammetric dense point clouds via bundle adjustment that capture the geometric structure of the plants. We provide detailed and accurate point-wise labels for individual plants, individual leaves, salient points on the leaves such as the tip and the base. Additionally we include measurements of phenotypic traits performed by experts from the German Federal Plant Variety Office (”Bundessortenamt) on the real plants, allowing to evaluate approaches not only on segmentation and keypoint detection, but also directly on the downstream tasks. The provided labeled point clouds enable fine-grained plant analysis and opens the door for further progress in the development of automatic phenotyping approaches, but also enable further research in closely related application areas such as surface reconstruction, point cloud completion, and semantic interpretation of point clouds.
I Introduction
Figure 1: Our dataset, called BonnBeetClouds3D consists of high-resolution point clouds extracted via bundle adjustment of high-resolution images captured from multiple viewing angles by a UAV. Together with the point clouds, we also provide accurate per-point labels of the plants, leaves, and salient points of the leaves, such as leaf tips, corners and bases. Furthermore, we provide ground truth values of measurements that are relevant for phenotyping, such as leaf lenghts and widths.
Agricultural production of food, feed, and fiber is facing severe challenges in the next decades. While the world population is steadily growing, inducing a higher demand for agricultural products, the availability of arable land is decreasing due to climate change [21]. At the same time, we have to reduce the effects of conventional agricultural practices on our environment calling for a more sustainable agricultural production in order to preserve biodiversity [25] and the productivity of arable lands [29, 19, 6]. As one measure, agricultural production needs to significantly reduce the amount of applied agrochemicals, such as fertilizer and herbicides.
A potential pathway to enable a more sustainable agricultural production is the breeding of more resilient crop varieties, and novel field management techniques including more targeted interventions at a plant level, and an increase of non-chemical weeding practices to reduce the amount of needed herbicides. All these areas benefit from artificial intelligence-based solutions, especially advances in robotics and computer vision. In particular, the development of more capable perception systems for robotic platforms or automatized tractors could lead to a higher automation enabling plant-specific treatments, but also novel ways of analyzing data from the field could lead to advancement towards more sustainable agricultural production.
In this paper, we target the field of plant trait assessment by observing the phenotype of a plant, which is an expression of its genotype, influenced by environmental conditions, and affected by the field management. The assessment of the morphological characteristics of a plant is to this date often performed by manual measurements or visual scoring of a trained expert. These experts evaluate different traits such as growth stage, leaf attitude, plant height and leaf length Fig. 2. However, this assessment is a tedious and labour-intensive task, which is furthermore subjective to a certain extent. All this affects the spatial extent, reliability, and repeatability of phenotyping activities.
Figure 2: The experts Facundo Ispizua Yamati and Dirk Koops estimating phenotypic traits in the fields.
A potential solution to the aforementioned limitations of conventional phenotyping is the usage of algorithmic solutions to automate the process, which would also lead to repeatable and objective measurements of phenotypic traits. In particular, a combination of a robotic solution that automatically acquires the relevant data in combination with a robot perception or computer vision-based approach to extract such traits from the recorded sensor data is a way to remove the so-called phenotyping bottleneck.
The significant progress in image-based perception, but also the development of increasingly capable perception systems in autonomous driving, has been propelled by the availability of domain-specific datasets [8, 11, 5, 1, 26, 2] to a large extent. While there has been recently an increasing interest in agricultural robotics, this progress is stifled by the availability of data to study and advance perception in this domain. There is an increasing number of datasets available in the agricultural domain for semantic interpretation of 2D field images [4, 23, 28], not many, however, contain 3D data which is highly relevant for morphological trait estimation [24].
With this dataset paper we focus on sugar beets, a key value crop in Germany, which is affected by a growing number of diseases, and therefore needs the development of robust varieties to avoid losses in productivity [20].
The main contribution of this paper is a large dataset covering 48 varieties of sugar beet plants, with point-wise annotations for individual plants and individual leaves enabling the development of segmentation algorithms. To allow for direct evaluation of downstream tasks such as extraction of phenotypic traits, we also provide reference values for commonly evaluated metrics such as leaf length, leaf width and stem length alongside the position of salient points such as the leaf tips, leaf corners and plant centers.
In sum we make the following contributions. Our dataset enables: (i) Studying phenotyping using photogrammetric point clouds, by exploiting geometric information; (ii) Instance segmentation of plant organs using 3D point clouds; (iii) Estimation of key locations on the leaves that are useful for phenotyping. These claims are backed up by the paper and our experimental evaluation.
II Related Work
There has been increasing interest in semantic interpretation of field images targeting semantic [14, 16, 15, 12, 13, 18], instance [27] but also panoptic segmentation [22] of crops and weeds. These approaches mainly use large image datasets with pixel-wise annotations of crops and weeds [4, 23], but also leaf annotations [10, 28]. For a broader overview of agricultural datasets for semantic interpretation, we refer to the survey by Lu et al. [17].
While there are several image datasets available, few datasets provide point cloud data revealing the geometric structure of the plants that seems mandatory for a fine-grained structural analysis of phenotypic traits. Chaudhury et al. [3] proposed a synthetic dataset generated based on plant models with the addition of noise. While being able to generate large amounts of data, the gap between simulation and reality poses a challenge for the deployment of approaches developed on this data to real fields. The Pheno4D dataset [24] provides point clouds of real tomato and maize plants at different growth stages captured with a high-precision laser scanner with accurate labels of individual plant organs. In contrast to our dataset, these point clouds were acquired in a lab setting under controlled conditions. Dutagaci et al. [7] instead provide a dataset with real world data, but it contains only rose plants. Khanna et al. [9] provide a dataset containing sugar beet plants, but they do not provide point wise instance labels nor reference measurements for leaf measures.
Our dataset covers real breeding trials showing real-world lighting conditions, but also occlusions induced by the planting of the crops. Furthermore, it is augmented with real measurements by the phenotyping experts from the german Federal Plant Variety Office (BSA) providing reference measurements of per-plot phenotypic traits.
Dataset Real data Field Semantic labels Instance labels Leaf measurements Varieties
Dutagaci [7] Rose
Khanna [9] Sugar beet
Pheno4D [24] Tomato, Maize
Ours Sugar beet
TABLE I: Performance of the baselines and our approach at filtering out bad predictions based on the confidence prediction.
III The BonnBeetClouds3D Dataset
Figure 3: Overview of the captured field. The data used for training of deep learning approaches is highlighted in purple, the validation area is highlighted in green and the test data is in the blue rectangle.
III-A Field Setup
We recorded the data on the sugar beet trial of the Federal Plant Variety Office in Magdeburg, Germany (N52.112028°, E011.558448°). They set up the trial in a single factorial block design with the factor variety which contained two repetitions. The plot size was 1.5 m by 7.0 m and contained three rows of plants with a row distance of $0.5$ m. We cover a total of 48 varieties in our dataset.
The field technicians of the BSA sowed the plants at a distance of $6$ cm inside the row and later manually thinned them out to a final distance of 18 cm by manual hoeing. They used plant protection products like herbicides in the normal course of business. Any remaining weeds were regularly removed by manual hoeing. In addition, they artificially irrigated the experiment when the weather conditions were too dry. They made these labor-intensive tasks to create ideal conditions for the development of the varieties.
III-B Sensor Setup
[…]
III-C Data Processing and Labeling
We process the recorded, overlapping 100-megapixel images of different view angles to determine the relative poses between the images via bundle adjustment, where we used the GPS positions as camera position initialization. Using the poses, we compute a photogrammetric point cloud of the complete field as shown in Fig. 3.
To obtain a reliably accurate segmentation of plants and leaves, we manually labeled the individual plants and leaf instances in the point clouds using the Semantic Segmentation Editor by Hitachi Automotive and Industrial Laboratory. See an example in Fig. 4.
Figure 4: In the top left part (a) we show an unlabeled point cloud with the original leaf colors, while in the bottom right (b) we show the labeled leaves.
Figure 5: The patch extraction process from the point cloud of the field. As can be seen in the field representation, a subset of plant has been annotated, here the leaf color encode the respective leaf instance. From the whole field point cloud we then extract patches, as the entire point cloud is too big for being processed at once. The colored dots represent the centers of all the patches extracted from that plot. The colored boxes, instead, show only a subset of the extracted patches for better visualization. On the right there are two example patches with color-coded leaf ids and grey for the unlabeled part.
The annotated plants were carefully checked by a person different from the one that annotated them in the first place, to ensure that multiple people checked each label in order to obtain a higher accuracy in the resulting annotations. There are many plants present for every variety, making it intractable and also unnecessary to label all of them. Therefore, we selected one patch of adjacent plants per variety and annotated them. We then produced patches of the point cloud that we extracted with a regular pattern over the breeding plots, and projected the labels into these patches. See an example in Fig. 5. We furthermore manually annotated different relevant keypoints of each individual leaf as shown in Fig. 6. The leaf base represents the point where the leaf blade joins the petiole, the leaf tip represent the point that is furthest from the petiole and the lateral corners represent the points of the leaf that are furthest away from the central axis of the leaf. The stem base, instead, is the point where the stem joins the beet crown, which we also define as the plant center.
Figure 6: Labeled keypoints. The leaf base is the point where the stem joins the leaf blade, the left and right corners are the two point furthest apart from the main leaf axis, which therefore can be used for the computation of the maximum leaf width, the leaf tip is the topmost point of the leaf and the stem base is the point where the stem joins the beet crown, which we also define as the plant center.
The BSA performed measurements of different variety traits of 30 leaves for every breeding plot. These are the leaf blade length and width, and the petiole length and width. Additionally, they provided us the average leaf angle with respect to the ground, average plant height, the curliness and the curviness of the leaf border for each plot.
III-D Data Organization
To ensure a fair evaluation of our approach we divided the dataset into a training set consisting of 1782 leaves from 128 plants to train deep learning approaches, a validation set of 260 leaves from 17 plants to tune the hyperparameter of the approaches and a test set of 619 leaves from 41 plants to validate the performance of the approaches. We provide labels for the train and validation sets. The point clouds are divided into patches of 1 m by 1 m by 1 m to make them usable directly for training deep neural networks. We use the ply format to deliver these patches, where each point has the attributes:
  • (x,y,z) coordinates
  • (nx,ny,nz) normals
  • (r,g,b) color channels
  • leaf_ids, a counter which starts from 0, -2 means unlabeled, -1 refers to the plant center
  • plant_ids, a counter which starts from 0, -1 means unlabeled
  • plot_id
  • keypoint_id, where 0 is the leaf base, 1 is the left corner, 2 is the leaf tip, 3 is the right corner and -1 is no keypoint
The keypoint ids are labeled as subsets of the leaves, with multiple points of the same leaf having the same keypoint id. To get the most precise keypoint position, we find the centroid all points of one leaf having the same keypoint id. The ground truth measurements of the the phenotypic traits are delivered for every plot in csv-format. These include:
  • leaf blade length
  • leaf blade width
  • petiole length
  • petiole width
IV Conclusion
In this paper, we presented a dataset which enables research on automated morphological parameter extraction for phenotyping. We provide reference values both for intermediate tasks like instance segmentation, and for downstream tasks like performing automated leaf measurements.
Acknowledgments
We thank the Bundessortenamt, especially Dr. Manthey, Dr. Lichthardt, Dr. Beukert, Dr. Borg, Ms. Zickenrott, and Mr. Bauske, for maintaining the experimental fields on which the data was collected, for supporting the data collection, for collecting the reference data for the phenotypic traits, for the technical advice on the current phenotyping practices, and for a pleasant collaboration.